OpenStudy (joftheworld):

the company also discovered that it costs$28 to produce 2 camera cases. $50 to produce 4 camera cases, and $140 to produces 10 camera cases.find the total cost of producing 12 camera cases.

OpenStudy (triciaal):

$28 for 2 cases or $140 for 10 cases means $14 per case cost of producing 12 = option 1 (10 + 2 ) = 140 + 28 = $168 option 2 12 * 14 = $168 $50 for 4 cases means 50/4 = 12.50 per case 12 cases at 12.50 each = $150 option 2 12/4* 50 = 3 * 50 = $150 "the company also discovered" means there is more to this problem and this will help to determine what rate to use to find the true cost of producing 12 cases do you have the rest of the problem?